From mar@roe.ac.uk Mon Dec 19 13:07:18 2011
Date: Mon, 19 Dec 2011 10:49:26 +0000
From: Mike Read <mar@roe.ac.uk>
To: VDFS Science Archive Team -- Clive Davenhall <acd@roe.ac.uk>,
     Eckhard Sutorius <etws@roe.ac.uk>, Mark Holliman <msh@roe.ac.uk>,
     Nicholas Cross <njc@roe.ac.uk>, Nigel Hambly <nch@roe.ac.uk>,
     Rob Blake <rpb@roe.ac.uk>, Ross Collins <rsc@roe.ac.uk>,
     Stelios Voutsinas <stv@roe.ac.uk>
Cc: CCs for VDFS -- Andrew Lawrence <al@roe.ac.uk>, Bob Mann <rgm@roe.ac.uk>,
     Jim Emerson <j.p.emerson@qmul.ac.uk>, Keith Noddle <keith@keithnoddle.org>,
     Lorenzo Rimoldini <Lorenzo.Rimoldini@unige.ch>,
     Mike Irwin <mike@ast.cam.ac.uk>, Norman Gray <norman@astro.gla.ac.uk>,
     Stephen Warren <s.j.warren@ic.ac.uk>, Tom Kerr <tkerr@jach.hawaii.edu>,
     Tom Shanks <tom.shanks@durham.ac.uk>
Subject: WFAU VDFS Science Archives team meeting minutes, 16/12/2011

Minutes of WFAU VDFS Science Archive meeting:  December 16th 2011
-------------------------------------------------------------------------
-------------------------------------------------------------------------

Present:   ETWS, RPB, NJC, ACD, MAR, MSH, STV
Apologies: NCH, JPE, KTN, AL, RSC, RGM, C37, 89X

DoNM: 10am Friday January 13th(!!) 2012 in the VH (NCH? in the chair)


Actions discharged this time:
-----------------------------

ACTION: Minute taker to make sure actions are fully explained with
         self-contained text descriptions. 
We'll see.


Actions partly discharged but continuing:
-----------------------------------------

ACTION: RPB to roll out latest curation client / DB server share
         configuration for routine ops use. 
After being initially confused by the over detailed action line, RPB 
decided this continues.

ACTION: RPB to try monthly detection table partitions for VVV.
         CU4 ingest of VVV data into separate partitions implemented,
         but further tests required 
Almost there. Later CUs not yet tested and assignment of multiframeID / 
partition for deep products not yet implemented.

ACTION: NJC/RSC to check how we are handling the replacement of deblend column
         with average confidence values in the standard catalogue output and
         update schemas as required.
Continues.

Actions carried forward from 2/12/11 meeting
---------------------------------------------

ACTION: RPB to set up WSA mirror/sync for flat file DBs using
         MS SQL replication.
The presence of mergelog tables in the VSA replication caused a clash 
with a survey release. So replication had been turned off! No-one could 
remember if mergelog tables needed to be in the sync DBs so they'll be 
removed and then see if that breaks anything. Continues

ACTION: ETWS, with RPB, to develop an automated helper script
         to keep release DB files links updated.
Continues

ACTION: RGM, with input from NCH, to contact VISTA PIs with suggested
         timescale for public releases.
Continues


Specific points and new actions:
--------------------------------

Project management:

The WFAU "Quarterly" meeting was held on the 7th December. Thanks to NJC 
for the mince pies.


WFCAM, VISTA and VST updates:

The VISTA lens top end was cleaned back in September resulting in a 0.1 
mag ZP improvement.


Comments and issues arising from CASU minutes:

The team noted the minutes of the meeting of 23/11/11; there were no 
comments.


Networking:

Up and running but ...
SRIF/UKLight would be down for maintenance for a short period on the 
morning of Tuesday 20th December.

WSA/VSA/OSA Operations:

P87 data (not VVV) have been CU4'ed.

Deep stacks are currently being produced, in parallel, for VVV P85.

ETWS reported:
-- Transferred and processed the EROS-2 revised catalogue.
-- Transferred the VISTA October data (P88).

ACD/ETWS had worked on and implemented the OSA schema in advance of the 
Durham ATLAS meeting.

ETWS had transferred a night's worth of ATLAS data and run CU3 against a 
test database.

The OSA browser and basic OSA web pages had been set-up together with 
test pre-release flat file listing and access.

The main points of note from the Durham Atlas meeting were that the 
illumination correction for VST is around 10% and more difficult to 
correct for; and there was a request for list driven (profile corrected) 
photometry for photo-z work.

Following on from work on the ERM for GES, ACD reported that he'd 
written schemas and produced/debugged a non MS SQL Server version of 
the database. A port to MS SQL Server would follow shortly.

Hardware and Systems:

The new Astrogrid/AstroDABis server is ready to host services.

The new cluster node is awaiting "electrification" or plugging in.

RGM is looking for a host server for the Monet DB tests, possibly use 
ramses10 but this is the public VVV server.


Software:

NJC reported that he had been working on the parallelisation of CU13 
code which is now employed in churning out VVV tile stacks. He had also 
fixed the CU16 issue that affected the production of GALEX neighbour 
tables and written some helper scripts to assist with the recent VMC and 
DXS product assignment problems.

ETWS reported:
-- Finished work on the broken monitor pages.
-- Work on fixing the CU4 memory problem continues.
-- Created schema for the EROS-2 source tables.


Survey Data Release:

A VIDEO database, VIDEOv20111208, had been released.

The later stage CUs for the next VMC release should be started week.

An issue with the DXS in DR9 had been reported by a user. This was down 
to an issue with the product assignment resulting in un-merged framesets 
and references to deprecated frames. NJC had tracked down and fixed most 
of the issues and a test release DB looks much better. The aim is to 
update the DXS in DR9 next week.


Non-survey Data Release:

Nothing to report this time.


Astrogrid deployment & Data Analysis services:

STV reported that he hoped to be able to start deploying some of the new 
UI against VVV TAP endpoints soon.

The new Astrogrid server would be taking over the Astrogrid services 
currently deployed at Leicester.

Miscellaneous:

Merry Christmas



-- 
Scanned by iCritical.